mathslover (mathslover):

prove centripetal force = m(v^2/r)

mathslover (mathslover):

\[\huge{m\frac{v^2}{r}}\]

mathslover (mathslover):

@UnkleRhaukus

mathslover (mathslover):

@waterineyes

mathslover (mathslover):

@Vaidehi09

OpenStudy (anonymous):

where is our starting point?

OpenStudy (anonymous):

i think we can start with v = rw?

mathslover (mathslover):

did u mean w--> omega ?

OpenStudy (anonymous):

yes

mathslover (mathslover):

K ! so f = ma a = v-u/t = rw-u/t

OpenStudy (anonymous):

If an object is undergoing uniform circular motion then \[R= r \cos \omega t \ i+ r \sin \omega t \ j\\ V=\frac{dR}{dt} = -r \omega \sin \omega t \ i+ r \omega \cos \omega t \ j \\ a=\frac{dV}{dt}= -r \omega^2 \cos \omega t \ i- r \omega^2 \sin \omega t \ j\\ \left| a \right|=r \omega^2\]

OpenStudy (anonymous):

then F=mrw^2 v=rw w=v/r -----> F=m v^2/r

OpenStudy (anonymous):

sorry, i lost my net there for a while. and @mathslover, don't proceed that way. @mukushla 's derivation is the right way.

mathslover (mathslover):

ok thanks a lot @mukushla and @Vaidehi09

OpenStudy (anonymous):

if u want @mathslover may i also contribute some

OpenStudy (anonymous):

When an object is in circular motion, it need centripetal force. A red ball is attached to a green cord (neglect its mass) passing through a small hole in a friction less, horizontal table. The red ball is initially orbiting in a circle of radius r with velocity v. A black ball is tied to the other end of the green cord. If it is in equilibrium, the gravitational force of the black ball Fg = Mg , provides the centripetal force Fc needed. \[Fc = m v^2/r = m w^2 r ( v=w*r )\] \[Fg = Fc \] then i.e. \[ Mg = m w^2 r \]

mathslover (mathslover):

thanks a lot to u too

OpenStudy (rishabh.mission):

Centripetal Acceleration, ac = v2/r ac = centripetal acceleration (m/s2) Centripetal Force According Newton's second law the centripetal force can be expressed as Fc = m ac = m v2/r (2) where Fc = centripetal force (N) m = mass (kg).

OpenStudy (anonymous):

\[F=mac\] where ac is centripetal acceleration therefor \[ac=v ^{2}/r\] \[F=mv^2/r\]

mathslover (mathslover):

thanks a lot @muhammad9t5and @rishabh.mission

mathslover (mathslover):

@muhammad9t5

OpenStudy (anonymous):

@mathslover friend then click on best answer to give me medal if you like. Thanks.

mathslover (mathslover):

I am also confused to give medal to whom !!!.. sorry